# Interpreting artificial neural networks to detect genome-wide association signals for complex traits

### v1

@@ Acknowledgements
BY, MA, LM were supported by the European Union’s Horizon 2020 Research and Innovation Programme under Grant agreements No 964874 Realment and 847776 CoMorMent. Thanks to the Agence Nationale de la Recherche (ANR-20-CE45-0010-01) for funding. Thanks to the High Performance Computing Center (HPC) of the University of Tartu for providing computational resources. Thanks to the Estonian Biobank Research Team (Andres Metspalu, Tõnu Esko, Reedik Mägi, Mari Nelis, Georgi Hudjashov) for the genotype and phenotype datasets used in this work. Thanks to Liis Karo-Astover for administrative support and proofreading. Data used for the gene-variant expression analyses described in this manuscript were obtained from the Genotype-Tissue Expression (GTEx) Portal on 04/04/2024.
@@
Conceptualisation: BY. Data curation: BY, MA. Methodology: BY. Investigation: BY, MA, FJ. Formal analysis: BY, MA. Visualisation: BY. Supervision: BY, LM. Original draft: BY. Review & editing: BY, MA, FJ, LM. Funding acquisition: LM.

### v2

@@ Acknowledgements
BY, MA, MNG, LM were supported by the European Union’s Horizon 2020 Research and Innovation Programme under Grant agreements No 964874 Realment and 847776 CoMorMent. Thanks to the Agence Nationale de la Recherche (ANR-20-CE45-0010-01) for funding. Thanks to the High Performance Computing Center (HPC) of the University of Tartu for providing computational resources. Thanks to the Estonian Biobank Research Team (Andres Metspalu, Tõnu Esko, Reedik Mägi, Mari Nelis, Georgi Hudjashov) for the genotype and phenotype datasets used in this work. Thanks to Liis Karo-Astover for administrative support and proofreading. Data used for the gene-variant expression analyses described in this manuscript were obtained from the Genotype-Tissue Expression (GTEx) Portal on 04/04/2024.
@@
Conceptualisation: BY. Data curation: BY, MA. Methodology: BY. Investigation: BY, MA, FJ. Formal analysis: BY, MA, MNG. Visualisation: BY. Supervision: BY, FJ, LM. Original draft: BY. Review & editing: BY, MA, MNG, FJ, LM. Funding acquisition: LM, FJ, BY.
@@
Supplementary Table Legends
S1 Table. Detected PAL including all SNPs with mapped genes.